• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Annand

    Log entry time 08:13:14 on March11,2005

    Entry number 141192

    keyword=Hall-A Owl Shift Summary

    Problems
    Shifted started with a power-supply problem on Q1.
    Controlled access necessary to reset the PSU.
    Q2 tripped off, but was reset using the standard
    GUI tool.
    BigBite QDCs giving errors. Reboot(s) of ROC5
    did not clear the problem. Controlled access necessary
    to power-cycle the crate....this cleared front-panel error lights

    Detector Configuration x>1 d
    LHRS 19.5 deg., P0=3.762 GeV
    BB -97.6 deg., P=0.324 GeV
    Neutron -97.6 deg., P=0.324 GeV

    Perform Harp scan, 5uA
    Set up 2x2 raster, 5uA on BeO target
    Switch to 4cm LD2 target
    Start scan of BB and Neutron rates vs beam current...
    5uA, 10uA, 15uA, 20uA
    Stop at 20uA...so not to exceed approx 1MHz on BB dE counters

    Run Summary
    2557 (?k events) Raster Check
    2558 (?k events) Raster Check
    2559 No record, ROC5 rebooted
    2560 (50k events) 5uA on LD2 (No BigBite QDC)
    2561 (50k events) 10uA on LD2 (No BigBite QDC)
    2562 (50k events) 15uA on LD2 (No BigBite QDC)
    2563 (50k events) 20uA on LD2 (No BigBite QDC)
    ROC5 rebooted
    2564 (?k events) 20uA on LD2 (No BigBite QDC)
    2565 (?k events) 20uA on LD2 (No BigBite QDC)
    ROC5 PSU cycled
    2566 (?k events) Check BigBite QDCs working
    2567 (50k events) 20uA on LD2 (BigBite OK)
    2568 (200k events) 20uA on LD2 (BigBite OK)
    2569 (200k events) 20uA on LD2 (BigBite OK)